# Two statistical problems for multivariate mixture distributions

### v1

@@
Abstract.
In an earlier work [12], it was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. Using this work, we address the following two important statistical problems: that of testing and measuring the agreement between two different random partitions, and that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based of univariate projections. We also compare our proposal with robust versions of the expectation-maximization method EM. In each case, we present algorithms for effecting the task, and compare them with existing methods by carrying out some simulations.
Key words and phrases:
@@ Abstract.
1. Introduction
Some recent results refine the classical Cramér–Wold theorem, and show that in some cases just finitely many projections suffice to determine the measure. See for instance [11] and the references therein, which include the case of Gaussian measures where just $(d^{2}+d)/2$ projections suffice.
Recently, in [12] these results have been extended to the case of mixtures, namely convex combinations of probability measures. In particular, the important cases of multivariate Gaussian and $t$-distributions mixtures are addressed in [12].
With these results in hand we then consider a new approach to two interesting statistical problems. The first one is a way to measure the agreement between two partitions of a data set, in particular, two outputs of different clustering procedures. The second one is to provide a new procedure for estimating a mixture of multivariate normal and $t$-distributions based on univariate projections.
@@ 1. Introduction
Section 2 begins with a short introduction to the notion of the Cramér–Wold device. In Section 3 we discuss the problem of comparing random partitions by model-based methods. Comparing partitions is an important problem in particular in cluster analysis, and a large amount of work on the subject introduced several different proposals. The most popular measures are given by the Rand index [26], the adjusted Rand index (Hubert and Arabie [18]), the Jaccard index, the Dunn index, the silhouette index, and the Xie–Beni index among others. See for instance the reviews by Arabie and Boorman [1], Fowlkes and Mallows [10] among others. However, most of these existing indices lack real mathematical analysis, and almost no information exists about their distribution, see [20].
@@
Two algorithms implementing these ideas are described, and then illustrated with some simulations. The simulations indicate a high degree of agreement between several well-known distance-measures and the one that we propose.
Finally, in Section 4, we present a new procedure for estimating mixtures of multivariate normal distributions, and mixtures of $t$-distributions based on one-dimensional projections of the mixture based on characteristic functions (ECF (see [33] and [37]), and using those univariate mixtures to provide consistent estimators of the parameters of the multivariate mixture.
2. Some preliminaries and recent results to be used
In this section, we introduce the fundamental concepts necessary for understanding our work. In particular, we clarify what we mean by projecting a measure onto a subspace and how distributions can be constructed through mixtures of Gaussian or Student distributions.
2.1. Projections
Given a Borel probability measure $P$ on $\mathbb{R}^{d}$ and a vector subspace $H$ of $\mathbb{R}^{d}$, we denote the projection of $P$ onto $H$ as $P_{H}$, specifically defined as the Borel probability measure on $H$ given by
$P_{H}(B):=P(\pi_{H}^{-1}(B)),$
where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
Let $\mathcal{P}$ denote a family of Borel probability measures on $\mathbb{R}^{d}$, and let $\mathcal{H}$ represent a family of vector subspaces of $\mathbb{R}^{d}$ (not necessarily of the same dimension). We say that $\mathcal{H}$ constitutes a Cramér–Wold system for $\mathcal{P}$ if, for every pair $P,Q\in\mathcal{P}$,
$P_{H}=Q_{H}\leavevmode\nobreak\ (\forall H\in\mathcal{H})\quad\Rightarrow\quad P=Q.$
In this context, the original Cramér–Wold theorem states that the family of all lines in $\mathbb{R}^{d}$ constitutes a Cramér–Wold system for the family of Borel probability measures on $\mathbb{R}^{d}$.
2.2. Mixtures
@@
(1) $f_{\mu,\sigma}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\Bigl{(}-\frac{(x-\mu)^{2}}{2\sigma^{2}}\Bigr{)}\quad(x\in\mathbb{R}).$
@@
(2) $\frac{1}{(2\pi\det(\Sigma))^{d/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
@@
[…]
2.3. A Cramér–Wold theorem for Gaussian mixtures and $t$-mixtures
In this section, we address the problem of testing for equality between two Gaussian mixtures by analyzing a finite number of projections. The basic theorem supporting this approach consists of two key components. The first is the abstract result presented in Theorem 4.1 in [12]. The second is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions), established in [11, Theorems 1 and 2], which we will now summarize.
@@ 2.3. A Cramér–Wold theorem for Gaussian mixtures and $t$-mixtures
We now recall the Cramér–Wold theorem for Gaussian mixtures [12, Theorem 4.1].
Theorem 2.1.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
We also recall an analogue of Theorem 2.1 for mixtures of multivariate $t$-distributions, given in [12, Theorem 4.2], thereby allowing heavy-tailed distributions.
Theorem 2.2.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
2.4. Strong sm-uniqueness sets
Theorems 2.1 and 2.2 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [12, Theorem 4.3], which also provides a realistic method for generating them.
Theorem 2.3.

### v3

@@
Abstract
We address two important statistical problems: that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based of univariate projections, and that of measuring the agreement between two different random partitions. The results are based on an earlier work of the authors, where it was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. We also compare our proposal with robust versions of the expectation-maximization method EM. In each case, we present algorithms for effecting the task, and compare them with existing methods by carrying out some simulations.
keywords:
@@ Abstract
pacs:
pacs:
JEL Classification]D8, H51
MSC Classification]35A01, 65L10, 65L12, 65L20, 65L70
1 Introduction
1.1 Random projections in statistics
Random-projection (RP) techniques have become a cornerstone in high-dimensional statistical analysis. Over the last two decades, the theoretical and applied developments of RP have flourished. Early applications emphasized computational efficiency and scalability, particularly for clustering [8], matrix approximations, [70, 41], and kernel learning [65]. The work presented in [1] introduced more efficient sparse and structured random matrices, which proved crucial for large-scale implementations.
The best-known method for applications involving random projections is the one based on the Johnson–Lindenstrauss lemma, which has numerous applications, especially in image processing. Although there are many other proposals based on RP, in what follows we will briefly describe the Johnson–Lindenstrauss approach below and then focus solely on methods based on extensions of the Cramér–Wold theorem. We will describe each of them in more detail below.
1.1.1 RP via the Johnson–Lindenstrauss lemma
This lemma can be stated as follows.
Theorem 1 (Johnson–Lindenstrauss [46]).
Let $\aleph_{n}$ be a set of $n$ points in $\mathbb{R}^{d}$, let $0<\epsilon<1$ and let $q>(8\log n)/\epsilon^{2}$. Then there exists a linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{q}$ such that, for all $u,v\in\aleph_{n}$,
$(1-\epsilon)\|u-v\|^{2}\leq\|f(u)-f(v)\|^{2}\leq(1+\epsilon)\|u-v\|^{2}.$ (1)
Thus any set of points in high-dimensional Euclidean space can be linearly embedded into a much lower-dimensional space such that pairwise distances are approximately preserved. This result lays the theoretical foundation for the widespread application of RP in modern data science and machine learning.
A natural question that arises is how to find the linear map $f$. Dasgupta and Gupta [23] propose a solution based on the following result, which allows us to find $f$ in randomized polynomial time.
Theorem 2 (Dasgupta–Gupta [23]).
Let $\aleph_{n}$ and $\epsilon$ be as in Theorem 1, and let $q>(4\log n)/(\epsilon^{2}/2-\epsilon^{3}/3)$. Define $f(u):=\sqrt{(d/q)}PUu$, where $U$ is a $d\times d$ unitary matrix whose columns are iid random vectors uniformly distributed on the unit sphere in $\mathbb{R}^{d}$, and $P:\mathbb{R}^{d}\to\mathbb{R}^{q}$ denotes the projection onto the first $q$ coordinates. Then
$\mathbb{P}\Bigl{(}(1-\epsilon)\|u-v\|^{2}\leq\|f(u)-f(v)\|^{2}\leq(1+\epsilon)\|u-v\|^{2})~\forall u,v\in\aleph_{n}\Bigr{)}\geq 1/n.$
Recent studies continue to build on the Johnson–Lindenstrauss lemma, investigating its application in modern contexts, for example, as a method for dimensionality reduction [22], particularly in machine-learning applications [30], in high-dimensional clustering [2], and nearest neighbor search [49]. These studies demonstrate how RP can preserve important properties in high-dimensional spaces while dramatically reducing the computational complexity of various algorithms.
A nice property of the inequality (1) is that it does not depend on the dimension $d$. However, if $\epsilon$ is small then $q$ becomes very large.
1.1.2 RP via the Cramér–Wold theorem
A more recent and statistically grounded direction emerged in the work of Cuesta, Fraiman and Ransford [18], based on extensions of the Cramér–Wold theorem.
In its original form, established in [15], the Cramér–Wold theorem says that a multivariate distribution is uniquely determined by its one-dimensional projections. Subsequent work of Rényi [67], Gilbert [38], Bélisle–Massé–Ransford [6] and Cuesta–Fraiman–Ransford [18] showed that, under appropriate assumptions on the growth of its moments, a multivariate distribution is determined by a ‘sufficiently rich’ set of lower-dimensional projections, thereby allowing it to be determined from random projections onto low-dimensional subspaces.
Here is one such result. In what follows, given a Borel probability measure $P$ on $\mathbb{R}^{d}$ and a vector subspace $H$ of $\mathbb{R}^{d}$, we denote by $P_{H}$ the projection of $P$ onto $H$, defined as the Borel probability measure on $H$ given by
$P_{H}(B):=P(\pi_{H}^{-1}(B)),$
where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$. Also, if $x\in\mathbb{R}^{d}$, then $\langle x\rangle$ denotes the vector subspace spanned by $x$.
Theorem 3 (Cuesta–Fraiman–Ransford [18, Corollary 3.2]).
Let $P$ and $Q$ be Borel probability measures on $\mathbb{R}^{d}$, where $d\geq 2$. Assume that:
(i) the absolute moments $m_{n}=\int\|x\|^{n}\,dP(x)$ are finite and satisfy the condition $\sum_{n=1}^{\infty}m_{n}^{-1/n}=\infty$;
(ii) the set $\{x\in\mathbb{R}^{d}:P_{\langle x\rangle}=Q_{\langle x\rangle}\}$ is of positive Lebesgue measure.
Then $P=Q.$
Cuesta–Fraiman–Ransford focused on using RP not only as a computational tool but as a vehicle for inferring robust statistical structure. In particular, they demonstrated how statistical procedures random-projection tests can preserve desirable properties such as consistency and robustness while remaining computationally feasible.
This line of work differs fundamentally from algorithmic RP approaches. Rather than using RP solely for dimensionality reduction, Cuesta, Fraiman and their collaborators exploit the probabilistic distribution of projections to characterize centrality, perform goodness-of-fit tests [16], depth measure [20], detect outliers [56], and hypothesis tests [36]. For example, in their framework, distributions are projected onto randomly chosen lines, and classical univariate techniques are applied. Repeating this process over multiple projections yields statistical conclusions that are consistent with high-dimensional structures.
These methods have inspired extensions in several applied domains. For instance, random-projection depth was later adapted to functional data analysis, see [21] and [19]. The article [60] proposes tests for functional data using projections, while [4] explores robust estimators via projection-pursuit methods. These studies reinforce the versatility of random projections as not only an algorithmic trick but a robust statistical principle. Moreover, this framework aligns conceptually with the work of [26], which emphasized the geometric underpinnings of high-dimensional phenomena, although the focus in [26] is often more combinatorial.
In contrast to mainstream applications of random projection, such as kernel approximation [65] or compressed sensing [11], the approach developed in [17, 18] is distinguished by its nonparametric nature and inferential focus. Rather than focusing on data transformations that preserve structural relationships for machine learning, the latter work is more oriented to preserving the geometric properties of the data for inferential purposes. This distinction in focus underscores the role of RP in nonparametric inference, as it provides an intuitive tool for deriving statistical conclusions from data without imposing stringent parametric assumptions.
More recently, there have been some further refinements of the Cramér–Wold theorem showing that, in some cases, just finitely many one-dimensional projections suffice to determine the measure. For example, the following result shows that, in the case of elliptical distributions (which includes Gaussian and $t$-distributions), just $(d^{2}+d)/2$ projections suffice.
Theorem 4 (Fraiman–Moreno–Ransford [35]).
Let $v_{1},\dots,v_{d}$ be linearly independent vectors in $\mathbb{R}^{d}$, and let $S=\{v_{j}+v_{k}:1\leq j\leq k\leq d\}$. If $P,Q$ are elliptical distributions in $\mathbb{R}^{d}$ and if $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
In [34] these results were extended to the case of certain multivariate mixtures. It was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. This last result is described in more detail in §2, and its applications form the subject of the rest of the paper.
1.1.3 On the pros and cons of RP-methods
Despite its many advantages, the RP methodology is not without its limitations. One notable drawback is that while it offers dimensionality reduction, it does so at the expense of certain properties of the data. Although the Johnson–Lindenstrauss lemma guarantees approximate distance preservation, it does not ensure exact preservation, especially when data points lie on complex, nonlinear manifolds. This means that RP may distort some aspects of the data distribution, especially for applications like clustering or classification, where fine-grained distinctions between data points are critical, see [8].
Moreover, while RP has been shown to be robust in certain high-dimensional settings, its performance can be sensitive to the type of distribution or structure inherent in the data. For example, in datasets with heavy tails or outliers, RP techniques may not preserve the underlying distribution as effectively as more robust methods like those based on depth functions, see [20]. Thus, while RP is often computationally efficient, it may sometimes fail to capture subtle patterns that are crucial in some statistical problems.
A significant challenge also arises from the inherent randomness of RP. This randomness introduces variability in the performance of algorithms across different projections. While multiple projections typically improve the consistency of results, determining the number of projections necessary for a reliable outcome is still an open question in many applications, see [18]. This issue has spurred research on the optimal number of projections required for specific tasks, particularly in the context of robust statistics and outlier detection.
Despite these challenges, the advantages of RP remain substantial. Its ability to reduce dimensionality while retaining much of the essential structure of the data makes it a powerful tool in large-scale statistical learning. Moreover, the simplicity of random projections allows for faster computations compared to other dimensionality reduction techniques, such as principal component analysis (PCA) or t-SNE, which can be computationally expensive in high-dimensional spaces, see [78].
As modern datasets continue to grow in complexity and dimensionality, the statistical insight provided by projection-based depth and testing becomes increasingly relevant. It offers a complement—and at times a counterpoint—to the more engineering-oriented uses of RP found in neural networks [14], sketching algorithms [50], or NLP embeddings [77]. The blend of theoretical rigor and practical robustness in the works of Cuesta, Fraiman, Ransford and their collaborators provides tools that are especially suitable for applied statisticians working under real-world constraints, such as noisy data or nonstandard distributions.
1.2 Roadmap of the rest of the article
In this article we consider two problems: that of estimating for mixtures of multivariate normal distributions and $t$-distributions based on univariate projections, and that of measuring the agreement between two random partitions (the output of two different model–based clustering procedures). To solve these problems we make use of some of the results given in [34].
Section 2 begins with a short introduction to the notion of the Cramér–Wold device, and the main notions and results to be used are described.
In Section 3, we present a new procedure for estimating mixtures of multivariate normal distributions, and mixtures of $t$-distributions based on one-dimensional projections of the mixture based on characteristic functions (ECF), see [76] and [82], and using those univariate mixtures to provide consistent estimators of the parameters of the multivariate mixture.
@@ 1.2 Roadmap of the rest of the article
In Section 4 we discuss the problem of comparing random partitions by model-based methods. Comparing partitions is an important problem in particular in cluster analysis, and a large amount of work on the subject introduced several different proposals. The most popular measures are given by the Rand index [66], the adjusted Rand index [45], the Jaccard index, the Dunn index, the silhouette index, and the Xie–Beni index among others. See for instance the reviews by Arabie and Boorman [3], Fowlkes and Mallows [33], among others.
@@
Algorithms implementing these ideas are described, and then illustrated with some simulations. The simulations indicate a high degree of agreement between several well-known distance-measures and the one that we propose.
2 Some preliminaries and recent results to be used
In this section, we introduce the fundamental concepts necessary for understanding our work. In particular, we clarify the notion of mixture, and describe a version of the Cramér–Wold theorem for mixtures of Gaussian or Student distributions.
2.1 Mixtures
Mixture models are fundamental tools in statistics and machine learning, providing a flexible framework for representing heterogeneous data as combinations of simpler component distributions. Since the pioneering work of Pearson [61], they have been extensively used in tasks such as clustering [74], density estimation [48], and the modeling of latent structures [59]. Finite mixture models have also found applications across a broad range of disciplines, including astronomy, biology, genetics, economics, social sciences, and engineering [52].
Estimation rates for finite mixture distributions, that is, for the mixture parameters, are also studied in [12, 44]. A more detailed review of the literature, including specific cases involving Gaussian or Student $t$ mixture distributions, can be found in [51].
@@ 2.1 Mixtures
$f_{\mu,\sigma}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\Bigl{(}-\frac{(x-\mu)^{2}}{2\sigma^{2}}\Bigr{)}\quad(x\in\mathbb{R}).$ (2)
@@
$\frac{1}{(2\pi\det(\Sigma))^{1/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$ (3)
@@
where $\nu$ is a positive integer, $\mu$ is a vector in $\mathbb{R}^{d}$, and where $\Sigma$ is a positive-definite $d\times d$ matrix. Once again, the constant $c_{\nu,\mu,\Sigma}$ is chosen to ensure that $\int_{\mathbb{R}^{d}}f_{\nu,\mu,\Sigma}(x)\,dx=1$.
2.2 A Cramér–Wold Theorem for Gaussian mixtures and $t$-mixtures
In this section, we address the problem for equality between two Gaussian mixtures by analyzing a finite number of projections. The basic Theorem supporting this approach consists of two key components. The first is the abstract result presented in Theorem 4.1 in [34]. The second is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions), established in [35, Theorems 1 and 2], which we will now summarize.
@@ 2.2 A Cramér–Wold Theorem for Gaussian mixtures and $t$-mixtures
We now recall the Cramér–Wold Theorem for Gaussian mixtures [34, Theorem 4.1].
Theorem 5.
Let $P$ and $Q$ be convex combinations of $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(2m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
We also recall an analogue of Theorem 5 for mixtures of multivariate $t$-distributions, given in [34, Theorem 4.2], thereby allowing heavy-tailed distributions.
Theorem 6.
Let $P$ and $Q$ be convex combinations of $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(2m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
2.3 Strong sm-uniqueness sets
Theorems 5 and 6 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [34, Theorem 4.3], which also provides a realistic method for generating them.
Theorem 7.